OpenStudy (anonymous):

If g(x) = x^2 - 3x - 8, find g(-6) a. 46 c. –206 b. 62 d. 10

OpenStudy (amistre64):

just let x = -6 ..... what is (-6)^2? what is -3(-6)? what is -8 ? what is the sum of them all?

OpenStudy (anonymous):

i just need the answer

OpenStudy (amistre64):

then yo need to go someplace else that just hands out answers.

OpenStudy (solomonzelman):

\[\huge\color{blue}{ g(x)=x^2-3x-8 }\] \[t o~~~solve~~~for~~~g(4)~.~~~~~~~\color{blue}{ g(4)=(4)^2-3(4)-8 }\]\[\color{blue}{ g(4)=(4)^2-3(4)-8 =16-12-8=-4 }~~~~~So,~~~~g(4)=-4\] can you similarly solve for g(-6) ( just like I showed on g(4) ) ?